OpenStudy (anonymous):

Please Help! A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 12(1.03)n Part A: When the scientist concluded his study, the height of the plant was approximately 16.13 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 3 to n = 10, and what does it represent? (4 points

OpenStudy (anonymous):

I apologize in advance if the connection times out. If you type a message and I don't reply, please wait. My net connection is horrible and it sometimes says I am typing but I'm not.

OpenStudy (whpalmer4):

Hi Antrod, what's your question? Where are you stuck?

OpenStudy (whpalmer4):

\[f(n)=12(1.03)^n\]is what the growth function really looks like, right?

OpenStudy (anonymous):

Yes. Yes. Thats what it looks like.

OpenStudy (anonymous):

Thank you so much for staying!

OpenStudy (whpalmer4):

Great. It wouldn't make sense as you typed it — could just have written \[f(n)=12.36n\]if it was all multiplication, and many growth formulas have an exponential form. So, now that we know what the formula is, what has you stuck?

OpenStudy (whpalmer4):

stupid OpenStudy formatting bug...

OpenStudy (anonymous):

The whole question. I know that is all starts from A. How do we answer A? If it says I'm typing after I send this, I'm not.

OpenStudy (whpalmer4):

Okay, so part A is just asking for a reasonable range of values (of \(n\)) to plot the function to include the point where \(f(n) = 16.13\).

OpenStudy (whpalmer4):

do you understand the concepts of domain and range of a function?

OpenStudy (anonymous):

Okay. So what d we have to do?

OpenStudy (whpalmer4):

I asked you a question...

OpenStudy (anonymous):

Still there?

OpenStudy (anonymous):

Sorry. The net messed up. I sent the message of what do we do and I didn't see your question. Answering it now.

OpenStudy (anonymous):

yes. x is the domain and y is the range.

OpenStudy (whpalmer4):

I am, still waiting for you to answer my question about whether or not you understand the concepts of domain and range of a function.

OpenStudy (whpalmer4):

okay, good. sometimes it is open study that puts the messages up out of order... so we know the function, and we know a value of \((x,y)\) of interest. Uh, we know a value of \(y\), we do not know the corresponding value of \(x\), yet. First step is to find the value of \(x\) that gives us that value of \(y\). How would you do that?

OpenStudy (anonymous):

I did. x is the domain and y is the range. when there is more than one matching domain its not a function.

OpenStudy (anonymous):

12 divided by both sides?

OpenStudy (whpalmer4):

\[f(n) = 12(1.03)^n\]\[f(n) = 16.13\]\[16.13=12(1.03)^n\]\[\frac{16.13}{12} = \frac{12(1.03)^n}{12}\]\[1.344\approx 1.03^n\] \[n\approx\]

OpenStudy (whpalmer4):

try a few values in your calculator, and find one that gives you something less than 1.344, and something greater.

OpenStudy (whpalmer4):

Or, you could use your knowledge of exponents or logs to find \(n\) directly.

OpenStudy (anonymous):

I think I figured it out. Thanks.

OpenStudy (anonymous):

Trying to give a medal to you.

OpenStudy (whpalmer4):

once you have figured out values that will let you plot a section of the function that shows it going through the value 16.13, that's your domain for part A. for part B, think of what the graph represents at \(n=0\). remember that \(n\) is the number of days of growth. for part C, the average rate of change of a function \(f(x)\) between points \(x=a\) and \(x=b\) is simply \[\frac{f(b)-f(a)}{b-a}\]